Find all real numbers x which satisfy the following equation: (x-6)^5 - 5(x-6)^3= 0
Thank you!

Answer by lwsshak3(11628)

Find all real numbers x which satisfy the following equation:
(x-6)^5 - 5(x-6)^3= 0
(x-6)^3(x-6)^2-5)=0
(x-6)^3=0
x=6 (multiplicity 3)
..
(x-6)^2-5=0
(x-6)^2=5
x-6=±√5
x=6±√5
x=6+√5
or
x=6-√5
solutions: x=6, 6, 6, 6-√5, 6+√5